OpenStudy (anonymous):

Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5.

OpenStudy (johnweldon1993):

Varying inversely would mean you use the equation \[y = \frac{ k }{ x }\] first...we want to solve for 'k' to use later....so you have y = -6 x = -2 plug those into that equation and solve for 'k'...what do you get?

OpenStudy (anonymous):

-2=k/-6

OpenStudy (johnweldon1993):

Well you actually switched the 'x' and 'y' it should be -6 = k/-2 but close.. so now..solve for 'k' by multiplying both sides of the equation by -2..what does 'k' equal?

OpenStudy (anonymous):

k=12 ?

OpenStudy (johnweldon1993):

right....that is what we get for 'k' Now...we have a second part of the question...solving for y when x = 5 to do this we use that same equation \[y = \frac{ k }{ x }\] you now know....k = 12 and x = 5 so plug those in..and solve for y...

OpenStudy (anonymous):

what is 12/5 ?

OpenStudy (johnweldon1993):

12 / 5 equates to 2.4 so y = 2.4 when x = 5

OpenStudy (anonymous):

thanks! can you help with this: Assume that y varies inversely with x. If y = 8.8 when x = 2.4, find x when y = 2.1. Round to the nearest tenth.

OpenStudy (johnweldon1993):

This would be the same thing... use the equation \[y = \frac{ k }{ x }\] input your 'y' and 'x' to solve for 'k' first....then use that 'k' and your given 'y' value to solve for 'x'

OpenStudy (anonymous):

ok thanks for the equation. that will help tons! :D

OpenStudy (johnweldon1993):

No problem... Remember that varying inversely \[y = \frac{ k }{ x }\] if it says it varies directly you would use \[y = k x\] hope that helps!